OpenStudy (anonymous):

Solve −4(x + 10) − 6 = −3(x − 2).

pooja195 (pooja195):

−4(x+10)−6=−3(x−2) (−4)(x)+(−4)(10)+−6=(−3)(x)+(−3)(−2)(Distribute) −4x+−40+−6=−3x+6 (−4x)+(−40+−6)=−3x+6(Combine Like Terms) −4x+−46=−3x+6 −4x−46=−3x+6 Add 3x to both sides. −4x−46+3x=−3x+6+3x −x−46=6 Add 46 to both sides. −x−46+46=6+46 −x=52 Divied both sides by -1

OpenStudy (anonymous):

so my answer is 25?

pooja195 (pooja195):

No.

OpenStudy (anonymous):

i dont understand because these are my answer choices 24 25 26 27

pooja195 (pooja195):

Did u type the problem wrong?

OpenStudy (anonymous):

i typed the answers wrong and got confused lol sorry